OpenStudy (anonymous):

2x^2+3y^2=6xy... Can you get the derivative (y') for me?

OpenStudy (anonymous):

Yeah Of course I can Just Sit and Watch : )

OpenStudy (anonymous):

come on now

OpenStudy (anonymous):

\[2x^2 + 3y^2 =6xy\] Differentiate with respect to x \[4x + 6y \frac{dy}{dx} = 6x \frac{dy}{dx} + 6y\]

OpenStudy (zion):

lana is online XD LOL

OpenStudy (lalaly):

yes i am :D

OpenStudy (anonymous):

Now Just some Algebra

OpenStudy (anonymous):

\[\frac{dy}{dx} =\frac{4x - 6y }{ 6x -6y}\] See that so Good

OpenStudy (anonymous):

Thank me by Hitting the Good Answer ; )

OpenStudy (anonymous):

you good! thanks!